OpenStudy (imtiaz7):

A current of 5A through a batter is maintained for 30s and in this time 600J of chemical energy is transformed into electric energy (a) What is the e.m.f of the battery? (b) How much electric power is available for heating and other uses? @needforspeed @xxAshxx

OpenStudy (mayankdevnani):

@Imtiaz7 i think emf of the battery is its potential difference.

OpenStudy (imtiaz7):

yes u r right

OpenStudy (mayankdevnani):

and we know that \[\huge W=P \times t\] plug these values in this equation, \[\huge 600=P \times 30\] so find P ? and \[\huge P=\frac{W}{t}=\frac{600}{30}\] so \[\huge P=20 W\]

OpenStudy (mayankdevnani):

understood? @Imtiaz7

OpenStudy (mayankdevnani):

so P= 20 W and this is our second part answer

OpenStudy (mayankdevnani):

now, for the first part question,

OpenStudy (imtiaz7):

@mayankdevnani yes u r going good

OpenStudy (mayankdevnani):

\[\huge e.m.f=V\] and \[\huge \color{red}{P}=V \times I\] so, \[\huge \color{red}{V}=\frac{P}{I}\] plug these values in this equation, \[\huge \color{red}{V}=\frac{20}{5}=4 \color{red}{V}\] and this is our first part answer

OpenStudy (imtiaz7):

absolutely right! and thanks @mayankdevnani

OpenStudy (mayankdevnani):

any time :) @Imtiaz7